Will sold 225 candy bars. He earned 1.00 for each almond bar he sold. He earned .75 for each carmel bar he sold. Will earned a total of 187.50. How many of each type of candy bar did he sell?
Answer by Maths68(1474)

Total Sale = 225 candy bars
Almond bar = x
Carmel bar = 225-x
Total earning = 187.50
Almond * cost + Carmel * cost = Total earning
x*1 + 0.75*(225-x)=187.50
x+168.75-0.75x=187.50
x-0.75x=187.50-168.75
0.25x=18.75
0.25x/0.25=18.75/0.25
x=75
Almond bar = x = 75 bars
Carmel bar = 225-x = 225-75 = 150 bars